# Sparse Imagination for Efficient Visual World Model Planning

### v1

@@ A.2 Main Method Implementation Details
Our model is built on a pre-trained DINO [3] ViT backbone which is frozen during training. Specifically, we utilize DINO-ViT-S/16 on 224x224 image, which creates a 14×14 spatial tokens per image. Visual tokens from context frames are flattened as a input to the world model. Actions and proprioceptive vectors are linearly projected on a dimension of 10 respectively and concatenated to every visual token along the feature dimension axis. The world model is a causal transformer decoder, which predicts the tokens of the next timestep. The transformer has 6 attention layers, 16 attention heads per layer, and 2048 embedding dimensions. The world model is trained with Adam optimizer at a learning rate of 5e-4 without scheduler for 100 epochs, batch size of 32, dropout probability of 0.1. All experiments were conducted using at most 4 Nvidia RTX 3090s.
MPC rollouts utilize the Cross-Entropy Method (CEM) with action distribution defined as a Gaussian distribution. CEM is performed with planning horizon of 5, optimizing by mean and variance of top-10 action sequences among 100 candidate sequences. CEM optimization is done for 10 iterations and returns the mean of the action distribution as the final action sequence, executing all the planned action sequence for every MPC step. MPC iterates until the task is completed or until it reaches a maximum limit of 10 iterations.
A.3 HSIC Analysis Details
@@
We measure statistical dependence between two random variables $X$ and $Y$ using the Hilbert–Schmidt Independence Criterion (HSIC) [30], defined via kernels $k$ and $\ell$ on their respective domains. Let $K$ and $L$ be the $n\times n$ Gram matrices with entries $K_{ij}=k(x_{i},x_{j})$ and $L_{ij}=\ell(y_{i},y_{j})$, and $H=I_{n}-\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$ the centering matrix. Then the empirical HSIC is
@@
A.4 Attentive Probing Details
@@
Attentive probing is a nonlinear probing method priorly used for downstream tasks since visual tokens do not satisfy linear separable guarantee. We use attentive probing to fairly compare predictive power of varying number of tokens. Instead of linear probing, attentive probing uses cross-attention layer to pool all input features to a single learnable query vector. After query residual connection, output is fed into 2-layer MLP with GeLU and LayerNorm, projecting to original input dimension. Then, final linear layer predicts the given probing target.
@@
A.5 Token Selection Methods Details
@@
LTRP (Learning to Rank Patches; [21]) utilizes a pre-trained MAE (Masked Autoencoder; [36]) to learn a ranking model that assesses the importance of individual patch tokens. With 90% of the patch tokens masked, LTRP compares the reconstruction results (using $L_{1}$ distance) when each remaining token is individually removed versus when it is retained. Tokens contributing more significantly to the reconstruction error are considered more important. The ranking model is trained using ListMLE [41], optimizing the order of the top 20 items within each list. For the ranking model, we employed a ViT-Small architecture, while the MAE utilized a ViT-Base checkpoint. The ranking model was trained on ImageNet-1K for 50 epochs using four NVIDIA RTX 3090 GPUs. For aspects not explicitly mentioned, training configurations adhered to the defaults outlined in the original paper. The LTRP code is available here [^2] https://github.com/irsLu/ltrp, and the MAE ViT-Base checkpoint is from here [^3] https://github.com/facebookresearch/mae.
@@
Similar to [23, 24, 25], we gauge token importance by examining the attention map between the CLS token and other patch tokens within a pre-trained DINO encoder. Tokens receiving higher attention weights from the CLS token are deemed most significant. We implement token dropping by retaining the top-K tokens sorted by their attention weights. Our experiments utilized the DINO-ViT-S/16 checkpoint [^4] https://github.com/facebookresearch/dino, specifically leveraging the attention map from the final transformer block of the DINO encoder.
@@
STAR (SynergisTic pAtch pRuning; [25]) determines a comprehensive importance score for each layer by combining two distinct components: an intra-layer importance score, derived from the attention weights between the CLS token and individual patch tokens, and an inter-layer importance score, quantified using Layer-wise Relevance Propagation (LRP). The inter-layer scores were pre-computed on the ImageNet-1K dataset, consistent with the methodology outlined in the original publication. Given that the original STAR implementation was based on DeiT, we modified their codebase to ensure compatibility with DINO. The aggregated score for layer $l$, denoted $S_{l}$, is formulated as $S_{l}=S_{l}^{\text{intra}^{\alpha}}\times S_{l}^{\text{inter}^{(1-\alpha)}}$. We set $\alpha=0.3$, which is one of the values proposed in the original paper. Analogous to the Attention-Encoder strategy, token dropping was realized by sorting tokens according to their scores from the final encoder layer in descending order and preserving only the top-K. The official code of STAR can be found here  [^5] https://github.com/LannWei/STAR_ICLR.
@@
Inspired by [28], we employ a modified version of Agglomerative Token Clustering (ATC) for our clustering approach. Necessary modifications were made as the original ATC was primarily designed for single-frame applications, such as classification or segmentation. Adapting it for goal-conditioned world model-based planning necessitated a modified mechanism to address multiple history frames, and to match clusters between the goal image features and predicted latent features.
@@
Appendix B Additional Experiments
B.1 Grouped Attention
@@
B.2 Planning Robustness to Noises
@@
B.3 Dropout vs. Clustering in Execution Times
@@ Appendix C Visual Demonstations
Figure 11: Rollout demonstration on Block Pushing environment of Full, CLS, and our methods (60% Drop).
Figure 12: Rollout demonstration on Wall environment of Full, CLS, and our methods (50% Drop).

### v2

@@ A.2 Main Method Implementation Details
Our model is built on a pre-trained DINO (Caron et al., 2021) ViT backbone which is frozen during training. Specifically, we utilize DINO-ViT-S/16 on 224x224 image, which creates a 14×14 spatial tokens per image. Visual tokens from context frames are flattened as a input to the world model. Actions and proprioceptive vectors are linearly projected on a dimension of 10 respectively and concatenated to every visual token along the feature dimension axis. The world model is a causal transformer decoder, which predicts the tokens of the next timestep. The transformer has 6 attention layers, 16 attention heads per layer, and 2048 embedding dimensions. The world model is trained with Adam optimizer at a learning rate of 5e-4 without scheduler for 100 epochs, batch size of 32, dropout probability of 0.1. All experiments were conducted using at most 4 Nvidia RTX 3090s.
A.3 Planning Details
MPC rollouts utilize the CEM with action distribution defined as a Gaussian distribution. CEM is performed with planning horizon of 5, optimizing by mean and variance of top-10 action sequences among 100 candidate sequences. CEM optimization is done for 10 iterations and returns the mean of the action distribution as the final action sequence, executing all the planned action sequence for every MPC step. MPC iterates until the task is completed or until it reaches a maximum limit of 10 iterations.
For the PointMaze, Wall, PushT, and Block Pushing environment, the MPC horizon is set to $H=5$. For each optimization step, we sample 100 candidate actions and select the top 10 based on the distance computed between the current and goal observations. Note that in sparse imagination scenarios, the distance is computed using only a subset of patches rather than the full patches. Each CEM optimization is performed for 10 steps. For the Wall and Block Pushing environments, the maximum number of MPC iterations is set to 10, while for Pointmaze and PushT, we set the maximum iterations to 15. For the Granular and Rope environments only employ the CEM approach for 10 steps.
A.4 HSIC Analysis Details
@@
We measure statistical dependence between two random variables $X$ and $Y$ using the Hilbert–Schmidt Independence Criterion (HSIC) (Gretton et al., 2005), defined via kernels $k$ and $\ell$ on their respective domains. Let $K$ and $L$ be the $n\times n$ Gram matrices with entries $K_{ij}=k(x_{i},x_{j})$ and $L_{ij}=\ell(y_{i},y_{j})$, and $H=I_{n}-\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$ the centering matrix. Then the empirical HSIC is
@@
A.5 Attentive Probing Details
@@
Attentive probing is a nonlinear probing method priorly used for downstream tasks since visual tokens do not satisfy linear separable guarantee (Bardes et al., 2024). We use attentive probing to fairly compare predictive power of varying number of tokens. Instead of linear probing, attentive probing uses cross-attention layer to pool all input features to a single learnable query vector. After query residual connection, output is fed into 2-layer MLP with GeLU and LayerNorm, projecting to original input dimension. Then, final linear layer predicts the given probing target.
@@
A.6 Token Selection Methods Details
LHS.
Latin Hypercube Sampling (LHS; (McKay et al., 2000)) is used as a simple uniform baseline that enforces spatial coverage. We treat the $H\times W$ patch grid as a 2D domain and partition it into $K$ strata corresponding to the desired number of tokens. At each step, we sample one patch from each stratum, which yields a set of retained tokens that are approximately uniformly distributed across the image.
@@ A.6 Token Selection Methods Details
LTRP (Learning to Rank Patches; Luo et al. (2024)) utilizes a pre-trained MAE (Masked Autoencoder; He et al. (2022)) to learn a ranking model that assesses the importance of individual patch tokens. With 90% of the patch tokens masked, LTRP compares the reconstruction results (using $L_{1}$ distance) when each remaining token is individually removed versus when it is retained. Tokens contributing more significantly to the reconstruction error are considered more important. The ranking model is trained using ListMLE (Xia et al., 2008), optimizing the order of the top 20 items within each list. For the ranking model, we employed a ViT-Small architecture, while the MAE utilized a ViT-Base checkpoint. The ranking model was trained on ImageNet-1K for 50 epochs using four NVIDIA RTX 3090 GPUs. For aspects not explicitly mentioned, training configurations adhered to the defaults outlined in the original paper. The LTRP code is available here [^4] https://github.com/irsLu/ltrp, and the MAE ViT-Base checkpoint is from here [^5] https://github.com/facebookresearch/mae.
@@
Similar to Liang et al. (2021); Wang et al. (2022); Zhang et al. (2024b), we gauge token importance by examining the attention map between the CLS token and other patch tokens within a pre-trained DINO encoder. Tokens receiving higher attention weights from the CLS token are deemed most significant. We implement token dropping by retaining the top-K tokens sorted by their attention weights. Our experiments utilized the DINO-ViT-S/16 checkpoint [^6] https://github.com/facebookresearch/dino, specifically leveraging the attention map from the final transformer block of the DINO encoder.
@@
STAR (SynergisTic pAtch pRuning; Zhang et al. (2024b)) determines a comprehensive importance score for each layer by combining two distinct components: an intra-layer importance score, derived from the attention weights between the CLS token and individual patch tokens, and an inter-layer importance score, quantified using Layer-wise Relevance Propagation (LRP). The inter-layer scores were pre-computed on the ImageNet-1K dataset, consistent with the methodology outlined in the original publication. Given that the original STAR implementation was based on DeiT, we modified their codebase to ensure compatibility with DINO. The aggregated score for layer $l$, denoted $S_{l}$, is formulated as $S_{l}=S_{l}^{\text{intra}^{\alpha}}\times S_{l}^{\text{inter}^{(1-\alpha)}}$. We set $\alpha=0.3$, which is one of the values proposed in the original paper. Analogous to the Attention-Encoder strategy, token dropping was realized by sorting tokens according to their scores from the final encoder layer in descending order and preserving only the top-K. The official code of STAR can be found here  [^7] https://github.com/LannWei/STAR_ICLR.
@@
Inspired by Haurum et al. (2024), we employ a modified version of Agglomerative Token Clustering (ATC) for our clustering approach. Necessary modifications were made as the original ATC was primarily designed for single-frame applications, such as classification or segmentation. Adapting it for goal-conditioned world model-based planning necessitated a modified mechanism to address multiple history frames, and to match clusters between the goal image features and predicted latent features.
@@
A.7 Cluster Retention Entropy Details
To quantitatively measure the spatial bias of different token sampling methods, we introduce Cluster Retention Entropy. This metric assesses how uniformly a sampling strategy distributes its kept tokens across the spatial layout of an image. The calculation proceeds as follows:
  1. 1.
    Spatial Clustering: First, we establish a set of spatial anchors. We take all visual patch tokens from a single, random image (e.g., Block Pushing) and cluster them into $K$ groups based on their spatial coordinates using the K-Means algorithm. This gives us $K$ spatial regions of interest.
  2. 2.
    Distribution of Retained Patches: For a single image and a given sampling method, we identify the subset of patches that are retained. We then count how many of these retained patches fall into each of the $K$ pre-defined spatial clusters. This count creates a discrete probability distribution over the clusters.
  3. 3.
    Entropy Calculation: Finally, we compute the Shannon entropy of this distribution. A higher entropy value indicates that the retained patches are spread more uniformly across the spatial clusters, signifying low spatial bias (i.e., good coverage). Conversely, a low entropy value indicates that the sampling method is heavily biased, concentrating its attention on only a few spatial regions.
In the analysis presented in the main text, we computed the mean entropy for each method on the Block Pushing dataset. To ensure a robust comparison, we averaged the results across a range of settings, varying the number of clusters $K\in\{2,4,8,12\}$ and the token drop ratio from 10% to 90%.
A.8 Real-World LeRobot Planning Details
To perform goal-conditioned planning in the real-world LeRobot environment, we used the goal image shown in Figure 19. The corresponding goal proprioceptive state, which represents the robot’s joint information, was set to [23.19,17.32,12.97,79.24,4.68,16.58][23.19,-17.32,12.97,79.24,-4.68,16.58] (PickPlace) and [31.44,0.68,16.94,67.90,26.77,0.61][-31.44,0.68,16.94,67.90,-26.77,0.61] (Drawer).
(a) PickPlace
(b) Drawer
Figure 10: The goal images used for planning in the real-world LeRobot environment.
Appendix B Additional Ablations and Experiments
B.1 Hyperparameter
CEM Optimization Steps.
We selected 10 CEM optimization steps for consistency with the DINO-WM baseline, ensuring fair comparison. Our ablation study in the PushT environment (60 episodes, various drop ratios) validated this choice: 5 steps yielded 49.3% success, while 10 steps achieved 51.7%. Increasing to 15 steps (50.7%) offered no significant benefit. Thus, 10 optimization steps represent a well-justified trade-off between planning quality and computational cost.
Longer Planning Horizon in MPC.
To further assess the robustness of our approach, we evaluated its performance under a more demanding, longer planning horizon. Specifically, we conducted experiments in the PushT environment with the planning horizon extended to H=7. Under this challenging condition, the Full-Patch baseline, which utilizes all visual tokens for planning, achieved a success rate of 48.3%. Our sparse imagination method remained highly competitive, achieving a 50.0% success rate with a 20% token drop and 46.7% with a 40% drop. This demonstrates that our framework’s performance does not degrade relative to the Full-Patch baseline even on longer-horizon tasks, confirming the robustness of planning with a sparse subset of tokens.
Frame Skips in World Model Training.
We performed an ablation on the frame skip value in the Pointmaze environment. We found that a frame skip of 1, which processes every frame, resulted in a low success rate of 36.7%, likely due to redundant observations. In contrast, increasing the frame skip to 5 significantly boosted performance to 85.8%. However, further increasing the frame skip to 10 yielded no additional benefit, with the success rate remaining at 85.5%. These results demonstrate a clear point of diminishing returns and justify our final choice of using a frame skip of 5, which provides an optimal balance between task performance and computational efficiency.
B.2 Grouped Attention
@@
B.3 Planning Robustness to Noises
@@
B.4 Dropout vs. Clustering in Execution Times
@@
B.5 Generalization to Different Planning Paradigms
To demonstrate that our sparse imagination framework is not limited to sampling-based planning algorithms like the CEM, we evaluate its compatibility with a gradient-based trajectory optimizer (MPC-GD). We replace the CEM planner with a gradient descent-based optimizer in the PointMaze and Wall environments, while keeping the randomized grouped attention and sparse token selection mechanisms identical to the main experiments.
Table 8 presents the planning success rates across varying dropout ratios. The results indicate that our method maintains high performance even when utilized with gradient-based optimization. Notably, with a 50% token drop, the success rates remain comparable to the Full-Patch baseline (e.g., 94.0% vs. 94.0% in PointMaze, and 88.0% vs. 88.0% in Wall). This suggests that random token sparsification preserves the optimization landscape sufficiently for gradient-based planners to converge to effective solutions, confirming the general applicability of our approach across different planning paradigms.
Table 8: Performance of Sparse Imagination using a Gradient-Descent based Planner (MPC-GD) on PointMaze and Wall environments (Mean Success Rate, %).
Drop Ratio PointMaze Wall
Full (0%) 94.0 88.0
10% 94.0 90.0
30% 98.0 92.0
50% 94.0 88.0
70% 90.0 82.0
90% 78.0 78.0
B.6 Exploring Uncertainty-Aware Adaptive Dropout
While our main experiments characterize performance as a function of fixed sparsity levels (demonstrating robustness up to approximately 50% dropout), we do not claim that a single fixed ratio is universally optimal. In scenarios where the optimal drop ratio is unknown or task complexity fluctuates, an online adaptive mechanism becomes desirable. Since our world model is explicitly trained to be robust under random sparsification via randomized grouped attention, the Sparse Imagination framework naturally supports dynamic adjustment of the dropout ratio at test time without retraining.
To explore this capability, we implement a simple uncertainty-aware adaptive planner. We utilize the variance of latent rollouts across MPC samples as a proxy for prediction uncertainty. The logic is straightforward: if uncertainty increases between iterations, the planner reduces the drop ratio (retaining more tokens); if uncertainty is low, the drop ratio is increased to save computation.
We evaluate two variants: (1) Ada-Rand, which updates the scalar drop ratio based on the change in variance and resamples a new random mask; and (2) Ada-Inc, which updates the mask incrementally, preferentially keeping high-variance tokens. Table 9 summarizes the results. In the Wall environment, Ada-Rand achieves a 95.0% success rate (matching the strong fixed 50% baseline) with an average drop ratio of 26.1%. In PushT, Ada-Rand reaches 68.3% success (matching the 10–50% fixed range). Crucially, these adaptive methods avoid the sharp performance collapses observed with overly aggressive fixed sparsity (e.g., >>70% in PushT) by automatically settling at a safe sparsity level. While they do not necessarily outperform the best-tuned fixed ratio, they demonstrate the potential of our framework to support online optimization when the optimal sparsity is unknown.
Table 9: Feasibility study of Uncertainty-Aware Adaptive Planning variants (Mean Success Rate, %) and the resulting Average Drop Ratio (%).
Wall PushT
Method Success Rate Avg. Drop Success Rate Avg. Drop
Ada-Rand 95.0 26.1 68.3 29.0
Ada-Inc 91.7 29.8 65.0 35.8
B.7 World Model Prediction Quality
To assess whether our training mechanism affects the intrinsic quality of the world model’s predictions, we train the world models with an optional decoder and measure the image reconstruction quality of the predicted features using the Learned Perceptual Image Patch Similarity (LPIPS) metric (Zhang et al., 2018) (lower is better). Table 10 compares the LPIPS scores of the standard Full (full attention) baseline against our model trained with Randomized Grouped Attention (“Drop”), where both models decode from full patches at evaluation time.
Across various environments, the LPIPS values for our method (Drop) remain very close to the Full-Patch baseline. Notably, in visually or physically complex tasks such as Granular and Rope, our method achieves slightly better reconstruction scores, potentially due to the regularization effect of the grouped attention strategy. These results indicate that our training mechanism does not degrade the world model’s prediction quality under full-patch evaluation while enabling robust performance under sparse imagination at planning time.
Table 10: Comparison of World Model Prediction Quality (LPIPS, lower is better).
Environment Full Drop (Ours)
PointMaze 0.00028 0.00031
Wall 0.00074 0.00078
Granular 0.05607 0.04739
Rope 0.03347 0.01906
PushT 0.00561 0.00644
B.8 Detailed Planning Time Analysis
To investigate the source of computational efficiency, we profile the latency of major components within the planning loop. Table 11 presents the fine-grained latency breakdown for the PushT environment.
The analysis confirms that the dominant acceleration stems from the World Model’s Self-Attention mechanism. As shown in the tables, Self-Attention latency drops dramatically (e.g., in PushT: $100\%\to 28.8\%$ at 50% drop), closely following the expected quadratic reduction. In contrast, the Projection and MLP layers scale linearly, while fixed costs remain stable.
Note: The latency values reported here were measured in a separate additional profiling session and may exhibit minor variances compared to the end-to-end timings in Table 2.
Table 11: Latency breakdown per planning iteration on PushT. Format: Time in seconds (Reduction % compared to Full baseline).
Variable Costs (World Model) Fixed Costs
Drop (%) Self-Attn Proj (QKV) MLP (FFN) Encoder Env Step Others
0 87.4 (-0.0%) 30.5 (-0.0%) 34.5 (-0.0%) 11.9 2.8 10.8
10 73.3 (-16.0%) 27.5 (-9.9%) 31.1 (-9.7%) 11.7 3.4 10.3
20 57.2 (-34.6%) 24.6 (-19.2%) 27.9 (-19.0%) 11.9 4.7 11.1
30 47.0 (-46.2%) 21.6 (-29.2%) 24.5 (-28.8%) 11.9 3.4 10.6
40 32.9 (-62.3%) 18.7 (-38.7%) 21.1 (-38.7%) 11.8 3.5 11.6
50 25.1 (-71.2%) 15.5 (-49.3%) 17.5 (-49.2%) 11.8 2.8 10.3
60 16.0 (-81.7%) 12.5 (-59.2%) 14.0 (-59.3%) 11.8 4.0 10.5
70 10.6 (-87.9%) 9.5 (-68.8%) 10.7 (-69.1%) 11.8 2.4 11.1
80 5.0 (-94.3%) 6.4 (-79.2%) 7.1 (-79.3%) 11.8 3.3 10.4
90 2.4 (-97.2%) 3.5 (-88.5%) 3.8 (-88.9%) 11.7 3.4 10.3
B.9 Additional Analysis on Blind Spots
To provide a quantitative characterization of the “blind spot” phenomenon visualized in Figures 9 and 20, we perform a controlled analysis on the PushT environment. At each MPC planning iteration, we track the ground-truth position of the agent (blue ball) for all candidate action sequences. We project the agent’s position into the image space. We approximate the blue ball as a circle whose diameter equals one patch size and use this circle to determine which patch tokens it overlaps. A blind-spot event is counted if all tokens overlapping with the agent are dropped by the sampling mask for a given timestep.
Table 12 presents the blind-spot occurrence rates evaluated at 50% and 60% dropout ratios (60 episodes). At 50% drop, Random sampling produces blind spots in only 3.10% of states, whereas baseline methods fail to capture the agent in 17–26% of cases. At 60% dropout, while failure rates increase due to higher sparsity, Random sampling (12.26%) still maintains a lower blind-spot rate compared to baselines (22–33%). These results quantitatively demonstrate that Random sampling consistently yields fewer blind-spot events compared to other selection strategies.
Table 12: Quantitative analysis of Blind Spot occurrence rates (%) in the PushT environment (60 episodes). Lower is better.
Method 50% Drop 60% Drop
Random (Ours) 3.10 12.26
STAR 17.74 22.87
LTRP 18.56 27.05
Attention-Encoder 20.99 33.04
Attention-WM 26.60 31.17
@@ Appendix C Visual Demonstations
C.1 Imagined Rollout by World Model
(a) LIBERO
(b) Meta-World
(c) LeRobot PickPlace
(d) LeRobot Drawer
Figure 14: Imagined rollout of the LIBERO, Meta-World, real-world LeRobot Pickplace and Drawer by world model. Top row: Ground truth. Bottom row: Imagined rollout by world model.
C.2 Executed Rollout during Planning
Figure 15: Executed rollout demonstration on real-world PickPlace task in LeRobot environment (50% Drop, VLA-only, and Full-Patch). A green border indicates a successful episode, while a red border indicates incompleteness. Both our method and the VLA-only baseline succeed in under 9 seconds, while the Full-Patch planner fails to do so in the same timeframe due to its significant planning latency.
Figure 16: Executed rollout demonstration on real-world Drawer task in LeRobot environment (50% Drop, VLA-only, and Full-Patch). A green border indicates a successful episode, while a red border indicates incompleteness. Both our method and the VLA-only baseline succeed in under 10 seconds, while the Full-Patch planner fails to do so in the same timeframe due to its significant planning latency.
Figure 17: Executed rollout demonstration on Block Pushing environment of Full, CLS, and our methods (60% Drop).
Figure 18: Executed rollout demonstration on Wall environment of Full, CLS, and our methods (50% Drop).
C.3 Additional Blind Spot Examples
(a) MoCo-v3
(b) DINO-v3
Figure 19: Visualization Examples of the Blind Spot Problem in Additional Visual Encoders (60% Drop in PushT).
(a) Random
(b) Importance-Based (Attention-Encoder)
Figure 20: Visualization Example of the Blind Spot Problem in Real-World LeRobot Environment.
Appendix D LLM Usage
LLMs were utilized solely for improving the clarity, grammar, and style of the paper and supplementary materials. No LLM was used for research ideation, experimental design, data generation, or result analysis. The authors take full responsibility for the content of this submission.